The diagonal of a rectangle measures 8sqrt(2) inches. If the width is 8 inches less than the length, then find the dimensions of the rectangle.
Answer by rothauserc(4718)

use the Pythagorean Theorem to solve
:
let x be the length of the rectangle
:
x^2 + (x-8)^2 = (8sqrt(2))^2
:
x^2 +x^2-16x+64 = 64*2
:
2x^2 -16x +64 = 128
:
divide both sides of = by 2
:
x^2 -8x +32 = 64
:
x^2 -8x -32 = 0
:
use quadratic formula
:
x = (-(-8) + sqrt((-8)^2 -4 * 1 * (-32)))/2(1) = 4 + 4 * sqrt(3) = 4(1 + sqrt(3))
x = 4(1 - sqrt(3))
:
we reject the negative value for x
:
****************************************************
length of rectangle = 4(1 + sqrt(3)) is approximately 10.93 inches
width is 10.93 - 8 = 2.93 inches
:
check the answer
:
(10.93)^2 + (2.93)^2 = (8*sqrt(2))^2
:
119.46 + 8.58 = 128
:
128.04 approximately equals 128
*****************************************************
: